# Temporal Sparse Autoencoders: Leveraging the Sequential Nature of Language for Interpretability

### v1

@@ 4.5 A Case Study in Using Temporal SAEs for Dataset Understanding and Control
Feature
Steering
Strength
Temporal SAE Matryoshka SAE
Medical case reports or concepts
We find a rare case of a 40-year-old man
with a history of chronic renal failure and
diabetes mellitus who presented with acute
onset of right upper quadrant pain. The
patient had a history of recurrent episodes
We find the effects of a magnetic field
on the dynamics of a two-dimensional
(2D) electron gas in a quantum well are
investigated by using the density matrix
formalism. The effect of an external
[Uncaptioned image][Uncaptioned image]
We find reported cases of the association of
a solitary renal cyst with an associated renal
lesion, and review the literature. The authors
report two cases of solitary renal cysts in
which the cyst was associated with a renal
We find 10-year-old children with a
history of recurrent urinary tract
infections (UTI) are more likely to have
an increased risk of developing chronic
kidney disease (CKD) than those
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
We find rare cases of the association of the
following three findings in a single case. We
present a case of a 42-year-old man with a
history of smoking for 30 pack-years who was
found to have the following four findings on
chest radiography: (1) a subpleural
We find betweenstory-1208252525252
5252525252525252525252525252525
2525252525252525252525252525252
52525252525252525252525
Book titles and authors
We find co-authored with Dr. John C. Maxwell,
The 21 Irrefutable Laws of Leadership is a
must read for anyone who wants to be a leader
in any capacity. In this book, John and I share
our experiences as leaders and how we have
We find, by the aid of a new method,
to determine the number of roots of
[…]
Let a and b be two numbers such that
a b; let f be any function which is
[Uncaptioned image][Uncaptioned image]
We find edited by the University of California
Press, 1986. The book is a collection of essays
on the history of the Jews in Russia and the
Soviet Union. It was written by the late
Professor $A.$ Goldberg, who was one of the
We find morire, to die; and morte, death,
[…]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
We find book reviews of the first two books in
the series, The Book of the Last Things and
The Book of the Last Things as Things. The
[…]
The first one was published by Ignatius Press in
[…]
a-b-j-f-l-w-c-c-e-c-c-a-$c-c-c-c-c-$
$c-c-c-c-c-c-c-c-c-$
Table 2: Examples of steering with medical and literature features. T-SAEs respond to high-level feature steering at various strengths and properly change the semantics of generation while retaining coherency. In contrast, Matryoshka SAEs require precise tuning of steering strength and fail catastrophically by repeating tokens due to the local nature of their features. Outputs are colored green if they achieve at least a score of 2 (out of 3) for both intervention success and coherence.
Dataset understanding.
While safety-tuning, often via reinforcement learning with human feedback (RLHF), is an essential part of the LLM alignment process, very little is known about the composition of popular preference datasets, as datasets do not come with explicit explanations or reasoning justifying the preferences, and human annotators are known to be noisy, inconsistent, or biased [45], allowing for the introduction of spurious correlations, outliers, and contradictions in the data. Furthermore, manual exploration and grading of datasets is difficult due to potential harms to graders, the scale of datasets, and the general lack of understanding of what to look when exploring the data. We present a case study on the utility of T-SAEs for helping to understand safety data at scale by discovering features that safety labellers prefer when annotating data, as well as potential unknown spurious correlations that may need to be addressed when training safety policy models. Using T-SAE decompositions, we analyze Anthropic’s Helpfulness Harmfulness RLHF dataset [44] to surface such problems in the data. Specifically, we compute the mean T-SAE feature activations for each candidate completion and inspect the difference between activations for the chosen and rejected pairs over the ‘harmless’ split of the dataset.
As expected, many of the features that activate highly for the rejected data are relevant to unsafe topics or intuitive human preferences. We observe features such as physical touch and intimacy, etiquette and social behavior guidelines, crime and malicious activities, and violent or aggressive behavior descriptions. These features clearly capture the semantics that we expect to find in an alignment dataset. However, a second class of features also appears with seemingly much less relevance to the task at hand, such as legal and formal language, quotes and speech, and transition words and phrases. At first glance, these features may seem like random noise artifacts or an issue with the T-SAEs themselves; however, upon further inspection, we find that they are the result of a consistent trend in the data. Frequently in the HH-RLHF dataset both options provided to the labeler are undesirable [46], generally falling into two cases: two harmful options or one harmful and one unhelpful response (where unhelpful may mean incoherent or irrelevant to the prompt). In the second case, raters consistently prefer the unhelpful response over the harmful one, leading to various undesirable spurious correlations. More specifically, the chosen response is often a single incoherent or irrelevant sentence that does not pertain to the request, whereas the rejected response is compliant with the request and will often include harmful references to online resources, quotes from prominent, divisive figures, and generally just lengthier responses. For example, when given the prompt "Human: So what do you think about the immagration problem", the chosen response is "Assistant: Immigraion is a new. I like to build things and help people" whereas the rejected response includes extensive, hallucinated quotes by John Horgan. This may lead them to prioritize noncompliance over quality, resulting in a higher prevalence of formal language, quotes, and longer responses in the rejected, unsafe responses, as picked up by the T-SAE with the aforementioned features. Following this, we hypothesize that some of the seemingly irrelevant features that activate more consistently for rejected, harmful responses are correlated with text length (e.g. transition words or long quotes). We ensure that these results are indeed factors of the data and not simply the SAE by conducting a two-sided t-test measuring the relationship between response length and human preference, finding that rejected responses are on average $29.7\%$ longer than chosen responses with statistical significance ($p=9\mathrm{e-}10$). We provide examples of max-activating samples for our relevant and spuriously correlated features and additional analysis of their correlation with response length in Appendix B.1.
This case study provides an initial signal that SAEs can be used to explore data at scale and surface behaviors that may not be expected beforehand. By decomposing alignment data with T-SAEs, we were able to identify unknown spurious correlations in the dataset and verify that these correlations were statistically significant. This discovery pipeline can be used to inform downstream dataset filtering and refinement, to apply corrective finetuning, or to aid in steering to improve model behavior.
Steering. In order to further validate the downstream utility of training SAEs with higher-level, semantic content, we evaluate whether T-SAEs provide greater utility for steering, as past literature has found that SAEs are not useful for inference-time intervention [17, 16, 47]. In particular, we follow the evaluation scheme from [16], which measures the tradeoff between steering success (do the outputs contain the feature that was intervened on) and steering coherence (are the outputs coherent), averaging across 30 different features (please see Appendix B.2 for more information). We use Llama3.3-70b to grade intervention success and coherence, with manual verification of grading. In Figure 5, we find that T-SAEs pareto-dominate existing SAEs, including the best SAE available for the same model on Neuronpedia [39].
Our results indicate that steering with high-level, semantic features results in more coherent and relevant generations than baselines. Intuitively, steering on low-, token-level features will result in a simple overwriting of the next-token generation objection, resulting in low- or token-level changes in outputs and the increased prevalence of specific tokens. In fact, a common failure mode when steering with state-of-the-art SAEs is token repetition at aggressive steering strengths (Table 2, right). In contrast, steering with high-level features changes the model’s encoding of the semantics of the task, guiding the model to generate text with the semantic content of the feature rather than simply encouraging token repetition (Table 2, left). As we train high-level T-SAE features to activate consistently over a sequence, these features are more amenable to steering, which generally involves exciting a feature over an entire sequence and is thus in-distribution for temporally consistent features. We additionally find that steering with high-level features is more forgiving: it properly impacts generation at lower strengths and continues to impact generation semantically up to high strengths without failing, whereas steering with baselines requires precise tuning to find a “sweet spot,” otherwise it will have no impact or will result in catastrophic failure.
Table 3: Ablation study of training configurations for Pythia-160m Temporal SAEs.
FVE ($\uparrow$)
Fraction
Alive
($\uparrow)$
Activation
Smoothness ($\downarrow$, H)
Semantics ($\uparrow$) Context ($\uparrow$) Syntax ($\uparrow$)
Random Contrast 0.0 -0.05 0.0 -0.02 +0.11 -0.10
@@
4.6 Ablation Studies
In the following ablation studies (Table 3), we explore the effect of varying components in our T-SAE training pipeline. All results are reported as the difference between the ablation and the normal T-SAE with the implementation described in Section 4.1. We conduct semantics, context, and syntax sparse probing evaluations on MMLU with $k=5$. First, we conduct a hyperparameter sweep on the high- and low-level feature percentages, varying the proportions to 10:90 and 50:50. As expected, if we increase the size of our high-level split, T-SAEs better recover semantics and context but perform worse on syntax. We next study the impact of contrasting with a token sampled randomly from any previous token in the context window, the trt-rth token with $r<25$ rather than from the t1t-1st token. When contrasting with a random token, we incorporate even longer-range dependencies into our temporal constraint; as a result of this, we see a large increase in performance on the context task and a large decrease in performance on the syntax task, but with minor change to semantic performance. Depending on the interpretability application, this behavior may be preferable to that of the T-SAEs trained on the immediate previous token, highlighting the need to carefully investigate inductive biases and data priors when building unsupervised concept discovery methods. Finally, we explore the impact of the contrastive component of our loss term, and consider the naive baseline of a sample-wise temporal similarity loss, $\ell_{i}=\alpha\|\mathbf{z}^{(i)}_{t}-\mathbf{z}^{(i)}_{t-1}\|_{2}^{2}$, averaged over a batch. This naive approach enforces less structure in the high-level feature space, resulting in worse semantic and contextual performance, but allowing for better performance on standard reconstruction metrics.
5 Discussion
@@
The efficacy of “bottom-up” interpretability methods, which aim to discover and represent concepts learned by large neural networks in an unsupervised fashion, has been a fiercely contested topic. In recent years, dictionary learning methods such as Sparse Autoencoders were hailed as a triumph in the interpretability community, showing promise in their ability to uncover unexpected and novel concepts, and providing a potential path forward for steering and control. However, as the excitement wore off, it became clear that SAEs suffer from a variety of issues, one of which being that they systematically fail to capture the rich conceptual information that drives linguistic understanding, instead exhibiting a bias towards shallow, syntactic features, such as ‘‘the phrase ‘The’ at the start of sentences.’’ This lack of deeper semantic concepts could potentially be attributed to the underlying LLMs themselves; maybe they don’t truly understand semantics or pragmatics, and it’s thus unrealistic to expect that interpretability methods would find them. But more likely than not, current concept discovery methods simply aren’t good enough to reveal the types of features we generally are interested in and that LLMs likely encode.
@@
In this work, we propose that this is due to a fundamental issue with how dictionary learning methods for LLMs are trained. Language itself has a rich, well-studied structure spanning syntax, semantics, and pragmatics; however, current unsupervised methods largely ignore this linguistic knowledge, leading to poor feature discovery that favors superficial patterns over meaningful concepts. We focus on a simple but important aspect of language: semantic content has long-range dependencies and tends to be smooth over a sequence, whereas syntactic information is much more local. We propose a novel loss function for training SAEs such that a subset of features behaves smoothly over time, better extracting semantic features from data. Through experiments, we demonstrate that the features learned by Temporal SAEs are more semantically structured, with minimal loss to reconstruction performance. We also present a case study demonstrating how this semantic information can be valuable in practical, real-world applications, such as finding safety vulnerabilities.
@@ 6 Acknowledgements
This work is supported by the National Science Foundation under grants CIF 2312667, FAI 2040880, and CIF 2231707, IIS-2008461, IIS-2040989, and IIS-2238714; an AI2050 Early Career Fellowship from Schmidt Sciences; and research awards from Google, OpenAI, the Harvard Data Science Initiative, and the Digital, Data, and Design (D3) Institute at Harvard. The views expressed are those of the authors and do not necessarily reflect the official policies or positions of the funding organizations. UB is funded by the Kempner Institute Graduate Research Fellowship. AO is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2140743.
@@
  • A Additional benchmark, probing, and TSNE results.
    • A.1 Probing and benchmark results across splits.
    • A.2 Probing results on more datasets and models.
    • A.3 TSNE visualizations of baseline SAEs.
  • B Additional case study details.
    • B.1 Alignment data case study.
    • B.2 Steering case study.
Appendix A Additional Results
A.1 Metrics across high and low splits
In Figure 6 and Table 4, we report probing and benchmark results across feature splits for Temporal SAEs and Matryoshka SAEs. We find that the high-level T-SAE feature space is smoother and contains more semantic and contextual information than its low-level. Additionally, we find syntactic information is equally spread between high- and low-level features due to the Matryoshka training setup (see discussion on disentanglement in Section 4.2). In contrast, Matryoshka SAEs are less smooth and place most syntactical information in the high-level split and are worse at semantics and context tasks across splits.
Table 4: SAE Benchmarks split across high and low splits, when applicable.
Activation
Smoothness
High Low
Pythia 160m
Temporal SAE 0.12 0.09 0.17
Matryoshka SAE 0.12 0.12 0.13
BatchTopKSAE 0.13 - -
Gemma 2-2b
Temporal SAE 0.13 0.10 0.15
Matryoshka SAE 0.14 0.15 0.12
BatchTopKSAE 0.13 - -
A.2 Probing Results on More Datasets
Figure 7: Accuracy of probes trained on SAE decompositions of Wikipedia and FineFineWeb for various SAEs trained on Gemma2-2b, as well as probes trained directly on model latents (darkorange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity (setup from Kantamneni et al. [40]).
Figure 8: Accuracy of probes trained on SAE decompositions of data from FineFineWeb (top), MMLU (middle), and Wikipedia (bottom) for Pythia-160m, as well as probes trained directly on model latents (darkorange), with semantic labels (right), contextual labels (middle), and syntactic labels (right) with varying levels of probe sparsity.
A.3 More TSNE Results

### v2

@@ 4.5 A Case Study in Using Temporal SAEs for Dataset Understanding and Control
Dataset understanding. We use T-SAEs to analyze Anthropic’s Helpfulness Harmfulness RLHF dataset (Bai et al., 2022) to discover features that safety labelers prefer when annotating data, as well as potential unknown spurious correlations that may need to be addressed when training safety policy models. For each candidate completion, we inspect the difference between the mean T-SAE feature activations for chosen and rejected responses over the ‘harmful/harmless’ split of the dataset, similar to previous work on using SAEs to understand human preferences (Movva et al., 2025).
As expected, many of the features that activate highly for the rejected data are relevant to unsafe topics or intuitive human preferences. We observe features such as “physical touch and intimacy”, “etiquette and social behavior guidelines”, “crime and malicious activities”, and “violent or aggressive behavior descriptions”. These features clearly capture the semantics that we expect to find in an alignment dataset. However, a second class of features also appears with seemingly much less relevance to the task at hand, such as “legal and formal language”, “quotes and speech”, and “transition words and phrases”. At first glance, these features may seem like random noise artifacts or an issue with the T-SAEs themselves; however, upon further inspection, we find that they are the result of a consistent trend in the data. Frequently in the HH-RLHF dataset both options provided to the labeler are undesirable [46], generally falling into two cases: two harmful options or one harmful and one unhelpful response (where unhelpful may mean incoherent or irrelevant to the prompt). In the second case, raters consistently prefer the unhelpful response over the harmful one, leading to various undesirable spurious correlations. More specifically, the chosen response is often a single incoherent or irrelevant sentence that does not pertain to the request, whereas the rejected response is compliant with the request and will often include harmful references to online resources, quotes from prominent, divisive figures, and generally just lengthier responses. This may lead them to prioritize noncompliance over quality, resulting in a higher prevalence of formal language, quotes, and longer responses in the rejected, unsafe responses, as picked up by the T-SAE with the aforementioned features (Appendix B.1). In contrast, existing SAEs highlight features that are generally uninformative and unrelated to the data, such as “specific bicycle components,” “terms related to data management”, and “references to ecosystem dynamics and environmental conditions” (Figure 5). The discrepancies between the safety concepts recovered by T-SAEs and Matryoshka SAEs highlight the utility of consistent, high-level features over noisy, local features for data aggregation.
Figure 5: Left. We use SAEs to understand user preferences in the HH-RLHF Bai et al. (2022) dataset. The Matryoshka SAE appears to find more random features, whereas the T-SAE is able to capture safety-relevant features. The T-SAE also finds a spurious length correlation in the dataset, where rejected completions are longer than chosen completions. Right. We find that high-level features Pareto dominate baselines in their ability to steer LLMs. Specifically, they are better at preserving coherence while successfully changing the semantics of model generation.
Steering.
In order to further validate the downstream utility of training SAEs with higher-level, semantic content, we evaluate whether T-SAEs improve steering performance, as past literature has found that SAEs are not useful for inference-time intervention (Wu et al., 2025; Bhalla et al., 2025; Arad et al., 2025). In particular, we follow the evaluation scheme from (Bhalla et al., 2025), which measures the tradeoff between steering success (do the outputs contain the feature that was intervened on) and steering coherence (are the outputs coherent), averaging across 30 different features (see Appendix B.2 for more information). We use Llama3.3-70b to grade intervention success and coherence, with manual verification of grading. In Figure 5, we find that T-SAEs pareto-dominate existing SAEs, including the best SAE available for the same model on Neuronpedia (Lin, 2023).
Our results indicate that steering with high-level, semantic features results in more coherent and relevant generations than baselines. Intuitively, steering on low-, token-level features will result in a simple overwriting of the next-token generation objection, resulting in low- or token-level changes in outputs and the increased prevalence of specific tokens. In fact, a common failure mode when steering with state-of-the-art SAEs is token repetition at aggressive steering strengths (Appendix B.2, Table 7, right). In contrast, steering with high-level features changes the model’s encoding of the semantics of the task, guiding the model to generate text with the semantic content of the feature rather than simply encouraging token repetition (Appendix B.2, Table 7, left). As we train high-level T-SAE features to activate consistently over a sequence, these features are more amenable to steering, which generally involves exciting a feature over an entire sequence and is thus in-distribution for temporally consistent features. We additionally find that steering with high-level features is more forgiving: it properly impacts generation at lower strengths and continues to impact generation semantically up to high strengths without failing, whereas steering with baselines requires precise tuning to find a “sweet spot,” otherwise it will have no impact or will result in catastrophic failure.
4.6 Ablation Studies
In Table 2, we explore the effect of varying components in our Temporal SAE training pipeline. All results are reported as the difference between the ablation and the normal Temporal SAE with the implementation described in Section 4.1. We conduct semantics, context, and syntax sparse probing evaluations on MMLU with $k=5$. First, we conduct a hyperparameter sweep on the high- and low-level feature percentages, varying the proportions to 10:90 and 50:50. As expected, if we increase the size of our high-level split, Temporal SAEs better recover semantics and context but perform worse on syntax. We next study the impact of contrasting with a token sampled randomly from any previous token in the context window, the trt-rth token with $r<25$ rather than from the t1t-1st token. When contrasting with a random token, we incorporate even longer-range dependencies into our temporal constraint; as a result of this, we see a large increase in performance on the context task and a large decrease in performance on the syntax task, but with minor change to semantic performance. Depending on the interpretability application, this behavior may be preferable to that of the Temporal SAEs trained on the immediate previous token, highlighting the need to carefully consider the features we hope to find when using unsupervised concept discovery methods. Finally, we explore the impact of the contrastive component of our loss term, and consider the naive baseline of a sample-wise temporal similarity loss, $\ell_{i}=\alpha\|\mathbf{z}^{(i)}_{t}-\mathbf{z}^{(i)}_{t-1}\|_{2}^{2}$, averaged over a batch. This naive approach enforces less structure in the high-level feature space, resulting in worse semantic and contextual performance but allowing for better performance on standard reconstruction metrics.
Table 2: Difference in performance between ablation and normal Pythia-160m Temporal SAEs.
FVE Fraction Alive Smoothness (High) Semantics Context Syntax
Random Contrastive 0.0 -0.05 0.0 -0.02 +0.11 -0.10
@@ 4.6 Ablation Studies
5 Discussion and Conclusions
@@
The efficacy of “bottom-up” interpretability methods, which aim to discover and represent concepts learned by large neural networks in an unsupervised fashion, has been a fiercely contested topic. In recent years, dictionary learning methods such as Sparse Autoencoders were hailed as a triumph in the interpretability community, showing promise in their ability to uncover unexpected and novel concepts, and providing a potential path forward for steering and control. However, as the excitement wore off, it became clear that SAEs suffer from a variety of issues, one of which being that they systematically fail to capture the rich conceptual information that drives linguistic understanding, instead exhibiting a bias towards shallow, syntactic features, such as “the phrase ‘The’ at the start of sentences.” This lack of deeper semantic concepts could potentially be attributed to the underlying LLMs themselves; maybe they don’t truly understand semantics or pragmatics, and it’s thus unrealistic to expect that interpretability methods would find them. But more likely than not, current concept discovery methods simply aren’t good enough to reveal the types of features we generally are interested in and that LLMs likely encode.
@@
In this work, we propose that this is due to a fundamental issue with how dictionary learning methods for LLMs are trained. Language itself has a rich, well-studied structure spanning syntax, semantics, and pragmatics; however, current unsupervised methods largely ignore this linguistic knowledge, leading to poor feature discovery that favors superficial patterns over meaningful concepts. We focus on a simple but important aspect of language: semantic content has long-range dependencies and tends to be smooth over a sequence, whereas syntactic information is much more local. We propose a novel loss function for training SAEs such that a subset of features behaves smoothly over time, better extracting semantic features from data. Through experiments, we demonstrate that the features learned by Temporal SAEs are more semantically structured, with minimal loss to reconstruction performance. We also present a case study demonstrating how this semantic information can be valuable in practical, real-world applications, such as finding safety vulnerabilities and steering.
@@ 6 Acknowledgements
This work is supported in part by the National Science Foundation under grants CIF 2312667, FAI 2040880, CIF 2231707, IIS-2008461, IIS-2040989, IIS-2238714, the AI2050 Early Career Fellowship by Schmidt Sciences, and faculty research awards from Google, OpenAI, Adobe, JPMorgan, Harvard Data Science Initiative, and the Digital, Data, and Design (Dˆ3) Institute at Harvard. UB is funded by the Kempner Institute Graduate Research Fellowship. AO is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2140743.
@@
  • A Limitations and Future Work
  • B Practical Applications of T-SAEs
    • B.1 Dataset Understanding of $H$-RLHF
    • B.2 Steering with T-SAEs and baselines
  • C Additional benchmark, probing, and TSNE results
    • C.1 Probing and benchmark results across splits
    • C.2 Baseline SAEs for Figure 4
    • C.3 Absorption Results
    • C.4 Scaling Study on Llama-3.1-8b-Instruct
    • C.5 Probing results on more datasets and models
    • C.6 TSNE visualizations of baseline SAEs
  • D Additional examples of syntactic features in public SAEs
Appendix A Limitations and Future Work
Limitations.
In order to compute the temporal contrastive loss term, each training data batch must include a second batch of tokens corresponding to the previous token’s activations, requiring smaller batch sizes for the same memory budget. Additionally, in this work we only explored splitting the feature space once into a single high-level and low-level feature space, but prior literature in sparse coding has shown that incorporating multiple scales of features can help to further disentangle sparse hierarchical signals and allow for more granularity on the temporal scale (Ong and Lustig, 2016). However, increasing the number of splits would further increase memory and computational complexity costs.
Future Work.
Building off of the above, one promising future direction is to explore multiple temporal hierarchies, which could correspond to book-, paragraph-, sentence-, and token-level information. By including more hierarchies, we expect to see even more disentanglement of semantic features at different scales. Given that T-SAEs allow for consistent tracking of features over sequences, another promising future direction may be to use the features learned by SAEs as “state” trackers, allowing for easy detection of significant changes in model behavior.
In this work, we used the traditional contrastive loss formulation, which involves computing cosine similarities and taking a softmax. However, these loss terms are generally used in traditional neural network representations spaces, which occupy all of $\mathbb{R}^{d}$ and can encourage contrastive pairs to not only be orthogonal but even have a negative cosine similarity with each other. In contrast, the SAE feature space is sparse and non-negative, living in a completely different space where cosine similarity is bounded between $[0,1]$. In this case, there may be alternative formulations of the contrastive loss which are more amenable to the geometry of the SAE feature space. For example, we can formulate linear approximation to the contrastive loss. Without loss of generality, denote $s_{ij}=s(z^{(i)}_{t},z^{(j)}_{t-1})$, and consider one row of the contrastive loss, which is a softmax:
$\log\left(\frac{e^{s_{ii}}}{\sum_{j}e^{s_{ij}}}\right)$ $=e^{s_{ii}}-\log(\sum_{j}e^{s_{ij}})$
We know the similarity scores are bounded between $[0,1]$ and given the sparsity of the space, we can expect the similarities between contrastive pairs to be much closer to 0. We use the approximations $e^{x}\approx 1+x$ and $\log(1+x)\approx x$ for small $x>0$.
$\approx s_{ii}-\log(\sum_{j}1+{s_{ij}})$
$=s_{ii}-\log(N+\sum_{j}(1+s_{ij}))$
$=s_{ii}-\log(N)-\log(1+\frac{1}{N}\sum_{j}(1+s_{ij}))$
$\approx s_{ii}-\log(N)-\frac{1}{N}\sum_{j}(1+s_{ij}).$
Then, discarding the constant $\log(N)$ term and summing over $i$, we arrive at a linearized contrastive loss after summing over $i$ and negating:
$\frac{1}{N}\sum_{i}s_{ii}+\frac{1}{N^{2}}\sum_{j}(1+s_{ij})$ (1)
This loss could potentially 1) be more amenable to the geometry of the SAE feature space and 2) be faster to optimize due to its linearity.